Geary Machine Shop is considering a four-year project to improve its production efficiency. Buying a new machine press for $547,200 is estimated to result in $182,400 in annual pretax cost savings. The press falls in the MACRS five-year class (MACRS Table), and it will have a salvage value at the end of the project of $79,800. The press also requires an initial investment in spare parts inventory of $22,800, along with an additional $3,420 in inventory for each succeeding year of the project.

  

Required :

If the shop's tax rate is 32 percent and its discount rate is 13 percent, what is the NPV for this project? (Do not round your intermediate calculations.)
Answer #1

Step 1 calculate the depreciation schedule:

Depreciation Schedule as Per MACRS 5 year rates
Year Opening Balance Investment Depreciation Closing Balance
0 547200 547200
1 547200 0.2*547200=109440 437760
2 437760 0.32*547200=175104 262656
3 262656 0.192*547200=105062.4 157593.6
4 157593.6 0.1152*547200=63037.44 94556.16
5 94556.16 0.1152*547200=63037.44 31518.72
6 31518.72 0.0576*547200=31518.72 0
  • Opening balance = previous year's closing balance
  • Closing balance = opening balance+investment-depreciation

Step 2 CF schdule and NPV calculation:

CF schedule
Year Remark 0 1 2 3 4
Pretax savings Given 182400 182400 182400 182400
Depreciation Calculated above 109440 175104 105062.4 63037.44
EBIT Pretax savings-Depreciation 72960 7296 77337.6 119362.56
Tax 32% x EBIT 23347.2 2334.72 24748.032 38196.0192
EAT EBIT-Tax 49612.8 4961.28 52589.568 81166.5408
Depreciation Added back as non cash expense 109440 175104 105062.4 63037.44
Initial investment Given -547200
WCINV Given -22800 -3420 -3420 -3420 -3420
WCINV Recovery As per convention 36480
Salvage Value Given 79800
Tax on profit 32% x (salvage value-book value at the end of year 4) -4426.848
CF EAT+Depreciation+Investment+WCINV+WCINV recovery+Salvage value+Tax on profit -570000 155632.8 176645.28 154231.968 252637.1328
Discount factor formula 1/1.13^0 1/1.13^1 1/1.13^2 1/1.13^3 1/1.13^4
Discount factor 1 0.884955752 0.783146683 0.693050162 0.613318728
DCF CF x Discount factor -570000 137728.1416 138339.1652 106890.4905 154947.0849
NPV = Sum of all Discounted CF -32095.11794

NPV = -32095.12, as it is negative, the project is not worth investing.